\begin{thebibliography}{99}


%\bibitem{JS_Mu90} M.~Muster, \textit{Computing certain invariants of topological spaces of dimension three}, Topology \textbf{32} (1990), 100--120.

%\bibitem{JS_Mu90a} M.~Muster, \textit{Computing other invariants of topological spaces of dimension three}, Topology \textbf{32} (1990), 120--140.

%%%%%%%Introduction and observations

%\bibitem{Langfellner2014} Langfellner, J., Gizon, L., Birch, A.~C.\ 2014.\ Time-distance helioseismology: A new averaging scheme for measuring flow vorticity.\ Astronomy and Astrophysics 570, A90. 


\bibitem{Basu2016}
S. Basu, \textit{Global seismology of the Sun},
  Living Reviews in Solar Physics, \textbf{13} (2016), 2.
  %, {doi:10.1007/s41116-016-0003-4}.

\bibitem{Gizon2010} L. Gizon, A.~C. Birch, H.~C. Spruit,   \textit{Local helioseismology: three-dimensional imaging of the solar interior},\ Annual Review of Astronomy and Astrophysics \textbf{48} (2010), 289--338. 


%\bibitem{Loeptien2014} L{\"o}ptien, B., Birch, A.~C., Gizon, L., Schou, J.\ 2014.\ Image compression in local helioseismology.\ Astronomy and Astrophysics 571, A42. 

\bibitem{Hanasoge2016} S. Hanasoge, L. Gizon, K.~R. Sreenivasan, \textit{Seismic sounding of convection in the Sun},  Annual Review of Fluid Mechanics \textbf{48} (2016), 191--217. 




\bibitem{Duvall}
T.~L. Duvall Jr.,
S. M. Jeffferies, J. W. Harvey, M. A. Pomerantz, \textit{Time-distance helioseismology}, Nature {\bf 362} (1993), 430--432.

%%%%%Normal modes
%\bibitem{Hanasoge2017} Hanasoge, S.~M., Woodard, M., Antia, H.~M., Gizon, L., Sreenivasan, K.~R.\  \textit{Sensitivity of Helioseismic Measurements of Normal-mode Coupling to Flows and Sound-speed Perturbations},   ArXiv e-prints arXiv:1705.08204 (2017).


%\bibitem{Papini2015} Papini, E., Birch, A.~C., Gizon, L., Hanasoge, S.~M.\ 2015.\ \textit{Simulating acoustic waves in spotted stars}, Astronomy and Astrophysics \textbf{577} (2015), A145. 

%\bibitem{Liang2017} Liang, Z.-C., Birch, A.~C., Duvall, T.~L., Jr., Gizon, L., Schou, J.\ 2017.\ Comparison of acoustic travel-time measurements of solar meridional circulation from SDO/HMI and SOHO/MDI.\ Astronomy and Astrophysics 601, A46. 



%%%% forward problem

\bibitem{Gizon2017} L. Gizon, H.  Barucq, M. Durufl{\'e}, C.~S. Hanson, M. Legu{\`e}be, A.~C. Birch, J. Chabassier, D. Fournier, T. Hohage, E. Papini,  \textit{Computational helioseismology in the frequency domain: acoustic waves in axisymmetric solar models with flows},  Astronomy and Astrophysics \textbf{600} (2017), A35. 

\bibitem{Barucq2017} H. Barucq, J. Chabassier, M. Durufl\'e, L. Gizon, M. Legu\`ebe,
\textit{Atmospheric radiation boundary conditions for the Helmholtz equation}, Mathematical Modelling and Numerical Analysis (2017), submitted.

\bibitem{Gizon2002} L. Gizon, A.~C. Birch,  \textit{Time-Distance Helioseismology: the forward problem for random distributed sources}, The Astrophysical Journal \textbf{571} (2002), 966--986. 

\bibitem{Boening2016} V.~G.~A. B{\"o}ning, M. Roth, W. Zima, A.~C. Birch,  L. Gizon, \textit{Sensitivity kernels for flows in time-distance helioseismology: extension to spherical geometry}, The Astrophysical Journal \textbf{824} (2016), 49. 


\bibitem{Montjoie}
M. Durufl\'e, 
\textit{Numerical integration and high-order finite element methods applied to time-harmonic Maxwell equations}, PhD Thesis (2006),  ENSTA ParisTech, France.
%\bibitem[Nagashima et al.(2017)]{2017A&A...599A.111N} Nagashima, K., Fournier, D., Birch, A.~C., Gizon, L.\ 2017.\ The amplitude of the cross-covariance function of solar oscillations as a diagnostic tool for wave attenuation and geometrical spreading.\ Astronomy and Astrophysics 599, A111. 


%\bibitem[Schunker et al.(2016)]{2016A&A...595A.107S} Schunker, H., Braun, D.~C., Birch, A.~C., Burston, R.~B., Gizon, L.\ 2016.\ SDO/HMI survey of emerging active regions for helioseismology.\ Astronomy and Astrophysics 595, A107. 

%%%%Inverse problem: noise

\bibitem{Gizon2004} L.~Gizon, A.~.C.~Birch, \textit{Time-distance helioseismology: noise estimation}, The Astrophysical Journal {\bf 614} (2004), 472--489. 

\bibitem{Fournier2014} D.~Fournier, L.~Gizon, T.~Hohage, A.~C.~Birch,  \textit{Generalization of the noise model for time-distance helioseismology}, Astronomy and Astrophysics \textbf{567} (2014), A137. 

%\bibitem{Couvidat2005} Couvidat, S., Gizon, L., Birch, A.~C., Larsen, R.~M., Kosovichev, A.~G.\ 2005.\ Time-Distance Helioseismology: Inversion of Noisy Correlated Data.\ The Astrophysical Journal Supplement Series 158, 217-229. 

%\bibitem{Jackiewicz2008} Jackiewicz, J., Gizon, L., Birch, A.~C.\ 2008.\ High-Resolution Mapping of Flows in the Solar Interior: Fully Consistent OLA Inversion of Helioseismic Travel Times.\ Solar Physics 251, 381-415. 

%\bibitem{Svanda2011} {\v S}vanda, M., Gizon, L., Hanasoge, S.~M., Ustyugov, S.~D.\ 2011.\ Validated helioseismic inversions for 3D vector flows.\ Astronomy and Astrophysics 530, A148. 

\bibitem{Kosovichev1996}
A. G. Kosovichev, \textit{Tomographic imaging of the Sun?s interior},  Astrophys. J. Lett. {\bf 461} (1996), L55.



\bibitem{Pijpers1992} F. P. Pijpers,  M. J. Thompson, 
\textit{Faster formulations of the optimally localized averages method for helioseismic inversion}, Astronomy and Astrophysics  {\bf 262} (1992), L33--L36.

%\bibitem{Pijpers1994} F. P. Pijpers,  M. J. Thompson, \textit{The SOLA method for helioseismic inversion}, Astronomy and Astrophysics  {\bf 281} (1994), 231--240



%\bibitem{Backus1967} G. E. Backus, J. F. Gilbert, \textit{Numerical applications of a formalism for geophysical inverse problems}, Geophys. J. R. Astr. Soc. {\bf 13} (1967), 247--276.

%\bibitem{Chavent1998} G. Chavent, \textit{Least-squares, sentinels and subtractive optimally localized average, in Equations aux d\'eriv\'ees partielles et applications}, 345--356,  Gauthier-Villars, Paris, 1998.

%\bibitem{Lions1988} J. L. Lions, \textit{Sur les sentinelles des syst\`emes distribu\'es. Le cas des conditions initiales incompl\`etes}, C. R. Acad. Sci. Paris S\'er. I Math. {\bf 307} (1988), 819--823.

\bibitem{Jackiewicz2012} J. Jackiewicz, A.~C. Birch, A.~C., Gizon, S.~M. Hanasoge, T. Hohage, J.-B. Ruffio, M.~{\v S}vanda. \textit{Multichannel three-dimensional SOLA inversion for local helioseismology}, Solar Physics \textbf{276} (2012), 19--33. 


\bibitem{Pinsker1980} M.~S. Pinsker, 
\textit{Optimal filtering of square integrable signals in Gaussian white noise}, Problems of Information Transmission {\bf 16} (1980), 120--133.

\bibitem{Fournier2016} D.~Fournier, L.~Gizon,  M.~Holzke, T.~Hohage.\  \textit{Pinsker estimators for local helioseismology: inversion of travel times for mass-conserving flows}, Inverse Problems \textbf{32} (2016), 105002. 

%\bibitem{Nussbaum1999} M. Nussbaum, \textit{Minimax risk: Pinsker bound},   Encyclopedia of Statistical Sciences, Update Volume {\bf 3}, 451--460 (S. Kotz, ed.), John Wiley, New York (1999) 
%Minimax Risk: Pinsker Bound, Michael Nussbaum, Cornell University, From: Encyclopedia of Statistical Sciences, Update Volume (S. Kotz, Ed.), 1999. Wiley, New York. We give an account of the Pinsker bound describing the exact asymptotics of the minimax risk in a class of nonparametric smoothing problems. The parameter spaces are Sobolev classes or ellipsoids, and the loss is of squared L2-type. The result from 1980 turned out to be a major step in the theory of nonparametric function estimation.

%\bibitem[Barekat et al.(2016)]{2016A&A...595A...8B} Barekat, A., Schou, J., Gizon, L.\ 2016.\ Solar-cycle variation of the rotational shear near the solar surface.\ Astronomy and Astrophysics 595, A8. 


%\bibitem[Birch et al.(2016)]{2016SciA....2E0557B} Birch, A.~C., Schunker, H., Braun, D.~C., Cameron, R., Gizon, L., Lo ptien, B., Rempel, M.\ 2016.\ A low upper limit on the subsurface rise speed of solar active regions.\ Science Advances 2, e1600557-e1600557. 




%\bibitem{Loeptien2015} L{\"o}ptien, B., and 12 colleagues 2015.\ Helioseismology with Solar Orbiter.\ Space Science Reviews 196, 251-283. 


%\bibitem{Langfellner2015} Langfellner, J., Gizon, L., Birch, A.~C.\ 2015.\ Spatially resolved vertical vorticity in solar supergranulation using helioseismology and local correlation tracking.\ Astronomy and Astrophysics 581, A67. 



%\bibitem{Barekat2014} Barekat, A., Schou, J., Gizon, L.\ 2014.\ The radial gradient of the near-surface shear layer of the Sun.\ Astronomy and Astrophysics 570, L12. 


%\bibitem[Nagashima et al.(2014)]{2014SoPh..289.3457N} Nagashima, K., L{\"o}ptien, B., Gizon, L., Birch, A.~C., Cameron, R., Couvidat, S., Danilovic, S., Fleck, B., Stein, R.\ 2014.\ Interpreting the Helioseismic and Magnetic Imager (HMI) Multi-Height Velocity Measurements.\ Solar Physics 289, 3457-3481. 



%\bibitem{Gizon2012} Gizon, L., Birch, A.~C.\ 2012.\ Helioseismology challenges models of solar convection.\ Proceedings of the National Academy of Science 109, 11896-11897. 


%\bibitem{Schunker2011} Schunker, H., Cameron, R.~H., Gizon, L., Moradi, H.\ 2011.\ Constructing and Characterising Solar Structure Models for Computational Helioseismology.\ Solar Physics 271, 1-26. 

%\bibitem[Cameron et al.(2011)]{2011ascl.soft05004C} Cameron, R., Gizon, L., Daiffallah, K.\ 2011.\ SLiM: A Code for the Simulation of Wave Propagation through an Inhomogeneous, Magnetised Solar Atmosphere.\ Astrophysics Source Code Library ascl:1105.004. 


%\bibitem[Moradi et al.(2010)]{2010SoPh..267....1M} Moradi, H., and 21 colleagues 2010.\ Modeling the Subsurface Structure of Sunspots.\ Solar Physics 267, 1-62. 




%\bibitem{Gizon2009} Gizon, L., and 14 colleagues 2009.\ Helioseismology of Sunspots: A Case Study of NOAA Region 9787.\ Space Science Reviews 144, 249-273. 




%\bibitem{Hanasoge2008} Hanasoge, S.~M., Birch, A.~C., Bogdan, T.~J., Gizon, L.\ 2008.\ f-Mode Interactions with Thin Flux Tubes: The Scattering Matrix.\ The Astrophysical Journal 680, 774-780. 


%\bibitem[Gizon et al.(2006)]{2006ApJ...643..549G} Gizon, L., Hanasoge, S.~M., Birch, A.~C.\ 2006.\ Scattering of Acoustic Waves by a Magnetic Cylinder: Accuracy of the Born Approximation.\ The Astrophysical Journal 643, 549-555. 


%\bibitem[Gizon and Birch(2005)]{2005LRSP....2....6G} Gizon, L., Birch, A.~C.\ 2005.\ Local Helioseismology.\ Living Reviews in Solar Physics 2, 6. 

\bibitem{Agaltsov2016} A.~Agaltsov, \textit{M\'ethodes de reconstruction pour des probl\`emes inverses pour des
\'equations de type Helmholtz}, PhD Thesis (2016), Universit\'e Paris-Saclay, France.

\bibitem{Novikov1988}
R.~G.~Novikov, \textit{Multidimensional inverse spectral problem for the equation $-\Delta \psi + (v(x) - Eu(x)) \psi = 0$}, Functional Analysis and Its Applications {\bf 22} (1988), 263--272.

\bibitem{Nachman1988}
A.~I.~Nachman, \textit{Reconstruction from boundary measurements}, Annals of Mathematics {\bf 128} (1988), 531--576.


\bibitem{Zotov2017}
	D.~I.~Zotov, A.~S.~Shurup, O.~D.~Rumyantseva, 
\textit{Vector field reconstruction of flows using the Novikov-Agaltsov functional algorithm and the additive correlation method},
Bulletin of the Russian Academy of Sciences: Physics {\bf 81} (2017), 101--105.

\bibitem{Shurup-priv}
A.~S.~Shurup, O.~D.~Rumyantseva, \textit{Joint reconstruction of sound speed, attenuation and currents by Novikov-Agaltsov functional algorithm},   Acoustical
Physics (2017), in press.

%holography

\bibitem{Cameron2008} R. Cameron, L. Gizon, T.~L. Duvall Jr.\  \textit{Helioseismology of sunspots: confronting observations with three-dimensional MHD simulations of wave propagation}, Solar Physics \textbf{251} (2008), 291--308. 

\bibitem{Hanasoge2010} S.~M.~Hanasoge, D.~Komatitsch, L.~Gizon,   \textit{An absorbing boundary formulation for the strati\-fied, linearized, ideal MHD equations based on an unsplit, convolutional perfectly matched layer}, Astronomy and Astrophysics \textbf{522} (2010), A87. 

\bibitem{Hanasoge2013} S.~M. Hanasoge, L. Gizon, G. Bal,   \textit{Propagation of seismic waves through a spatio-temporally fluctuating medium: homogenization}, The Astrophysical Journal \textbf{773} (2013), 101. 

\bibitem{Hanasoge2011} Hanasoge, S.~M., Birch, A., Gizon, L., Tromp, J.\  \textit{The Adjoint Method Applied to Time-distance Helioseismology}, The Astrophysical Journal \textbf{738} (2011), 100. 


\bibitem{ref2} T. Hohage, S. Langer, \textit{Acceleration techniques for regularized Newton methods applied to electromagnetic inverse medium scattering problems}, {Inverse Problems}, {\bf 26} (2010),    {074011}.


\bibitem{Lindsey2000a} C. Lindsey, C., D.~C. Braun,   \textit{Basic Principles of Solar Acoustic Holography},  Solar Physics {\bf 192} (2000), 261--284. 

\bibitem{Lindsey2000b} C. Lindsey, D.~C. Braun,   \textit{Seismic Images of the far side of the Sun}, Science {\bf 287} (2000), 1799--1801. 

\bibitem{Skartlien2002}
R. Skartlien, Local helioseismology as an inverse source-inverse scattering problem, The Astrophysical Journal {\bf 565} (2002), 1348--1365.

%\bibitem{Lindsey1997} C. Lindsey, D.~C. Braun,  \textit{Helioseismic holography},  The Astrophysical Journal {\bf 485} (1997), 895--903. 

%\bibitem{Braun1992} D.~C. Braun, C. Lindsey, Y. Fan,  S.~M. Jefferies,  \textit{Local acoustic diagnostics of the solar interior}, The Astrophysical Journal {\bf 392} (1992), 739--745. 

%\bibitem{Porter1969}R.~P.~Porter, \textit{Image formation with arbitrary holographic type surfaces}, Physics Letters A {\bf 29} (1969), 193--194.

%\bibitem{Porter1970} R.~P.~Porter, \textit{Diffraction-limited scalar image formation with holograms of arbitrary shape}, Journal of the Optical Society of America {\bf 60} (1970), 1051--1059.

%\bibitem{Bojarski1973} N.~N.~Bojarski, \textit{Inverse scattering}, Naval Air Systems Command Report (1973), Section 11, 3--6.

\bibitem{Porter1982} 
R. P. Porter and A. J. Devaney,
\textit{Holography and the inverse source problem}, Journal of the Optical Society of America {\bf 72} (1982), 327--330.

\bibitem{Devaney1985}
A.~J.~Devaney, R.~P.~Porter,  \textit{Holography and the inverse source problem. Part II: Inhomogeneous media}, 
Journal of the Optical Society of America A {\bf 2} (1985), 2006--2012. 

%\bibitem{Tsang1987} L. Tsang, A. Ishimaru, R.~P. Porter, D. Rouseff, \textit{Holography and the inverse source problem. III. Inhomogeneous attenuative media}, Journal of the Optical Society of America A {\bf 4} (1987),  1783--1787.


%\bibitem{Porter1986} R.~P.~Porter, Scattered wave inversion for arbitrary receiver geometry, The Journal of the Acoustical Society of America {\bf 80} (1986), 1220.


%%%Outstanding problems

%%%%sunspots

%\bibitem{Liang2013} Liang, Z.-C., Gizon, L., Schunker, H., Philippe, T.\ 2013.\ Helioseismology of sunspots: defocusing, folding, and healing of wavefronts.\ Astronomy and Astrophysics 558, A129. 

%\bibitem{Schunker2013} H. Schunker, L. Gizon, R.~C. Cameron, A.~C. Birch, \textit{Helioseismology of sunspots: how sensitive are travel times to the Wilson depression and to the subsurface magnetic field?}, Astronomy and Astrophysics \textbf{558} (2013), A130. 

%\bibitem{Cameron2011} Cameron, R.~H., Gizon, L., Schunker, H., Pietarila, A.\ 2011.\ Constructing Semi-Empirical Sunspot Models for Helioseismology.\ Solar Physics 268, 293-308. 



%%%%Homogenization: random flows



\end{thebibliography}

\documentclass{owrart}

\usepackage{xcolor}

\newcommand{\bu}{\mathbf{u}}

\newcommand{\br}{\mathbf{r}}

\begin{document}


\begin{talk}[Damien Fournier, Thorsten Hohage]{Laurent Gizon}
{Problems in computational helioseismology}
{Gizon, Laurent}

\noindent

The Sun supports acoustic oscillations continuously excited by near-surface turbulent convection.  {\it Global helioseismology} consists of inverting the measured frequencies of the normal modes of oscillation to infer the sound speed and rotation as a function of radius and unsigned latitude \cite{Basu2016}. 
 Techniques of {\it local helioseismology} based on correlations of  the wave field at the surface are being developed to infer the structure and dynamics of the Sun in three dimensions  \cite{Gizon2010, Hanasoge2016}. 

{\bf Forward problem.} Time-distance helioseismology \cite{Duvall} is a particular technique of local helioseismology,  analogous to geophysical seismic interferometry. Ignoring terms that involve gravity, the oscillations at position $\br$ and frequency $\omega$ can be described by a scalar field $\psi(\br, \omega)$, which solves the acoustic wave equation  \cite{Gizon2017}
\begin{equation}
L_{\br,\omega} [\psi] : = 
-(\omega^2 + 2i \omega\gamma)\psi  -2i \omega \bu\cdot\nabla_{\br}  \psi - c \nabla_\br\cdot\left(\frac{1}{\rho } \nabla_\br ( \rho c \psi ) \right) = s(\br,\omega)   ,
\end{equation}
where $\gamma(\br,\omega)$ is attenuation and the steady background medium is represented by density $\rho(\br)$, sound speed $c(\br)$, and flow $\bu(\br)$. Waves are excited by a stationary random process (granulation) represented by the function $s(\br,\omega)$.
The above equation is supplemented by a radiative boundary condition  \cite{Barucq2017}.
The basic input data  in time-distance helioseismology is the covariance function $C(\br',\br,\omega)= \psi^*(\br', \omega) \psi(\br, \omega)$ between two points on the solar surface. Under the assumption that  sources are spatially uncorrelated and of the form $\mathbb{E}[s^*(\br',\omega) s(\br,\omega)] = \delta(\br-\br')   P(\omega) \gamma(\br,\omega)  / \rho(\br) $ we have (to within a surface term)  
\begin{equation}
 C (\br',\br, \omega)  = \frac{P(\omega)}{4i\omega} \left[  G (\br, \br', \omega)  - G^\dagger (\br, \br',\omega) \right]  +  \text{noise}  , 
 \label{eq:CGG}
 \end{equation}
where  $   L_{\br,\omega}[ G(\br,\br',\omega) ] =  \delta(\br-\br')/ \rho (\br) $ and 
$G^\dagger = G^*(\bu\rightarrow -\bu)$ is obtained by switching the sign of $\bu$ and taking the complex conjugate. 
The linear forward problem consists in computing the perturbations to the covariance function caused by infinitesimally small  perturbations in the background medium. Combining the first Born approximation \cite{Gizon2002, Boening2016}
and Eq.~(\ref{eq:CGG}), Gizon et al.  \cite{Gizon2017} expressed sensitivity kernels in terms of only four Green's functions in the reference medium,  computed using the finite-element code Montjoie \cite{Montjoie}.

{\bf Inverse problem.} 
The  inverse problem consists of reconstructing $\gamma(\br,\omega)$, $c(\br)$, $\rho(\br)$, and $\bu(\br)$ in the interior, starting from a reference solar model. This requires knowledge of the noise covariance matrix \cite{Gizon2004,Fournier2014}.
Linear inversions are traditionally performed using Tikhonov regularization \cite{Kosovichev1996} or the
method of approximate inverse (called optimally localized averaging, see ref.~\cite{Pijpers1992}). 
Under the assumption of local horizontal translation invariance of the sensitivity kernels, multichannel inversions in Fourier space enable to solve problems that would otherwise require too much computer memory \cite{Jackiewicz2012}. Minimax estimators have been computed for this problem using the Pinsker method \cite{Pinsker1980, Fournier2016}.

The non-linear inverse problem of time-distance helioseismology (finite pertubations to the medium) has not been studied in full detail yet. Future studies should build on existing theoretical uniqueness results, in particular on the  Novikov-Agaltsov reconstruction algorithm \cite{Agaltsov2016}, which combines measurements of $G$ at several frequencies (see table below).  For measurements of $C$ instead of $G$, we have conducted numerical experiments to determine the number of frequencies required to reconstruct $\rho$ and $c$. For realistic noise levels, more frequencies will be needed to obtain useful reconstructions.
\begin{table}[h]
\label{table:GFH}
\caption{Number of frequencies needed for reconstruction} 
\begin{tabular}{c | c c c c} 
\hline\hline 
  & \multicolumn{2}{c}{Observable: $G$} &
\multicolumn{2}{c}{Observable: $C$}
     \\ [0.5ex] %
 & theory  & experiment  & theory  & experiment \\ [0.5ex] %
 \hline 
$c$ & 1 (ref. \cite{Novikov1988})& 1 (ref. \cite{Nachman1988})& ? & 2 (this work)\\ 
$\rho$ &1 (ref. \cite{Novikov1988})& 1 (ref. \cite{Nachman1988})& ? & 2 (this work)\\
$c$, $\rho$  & 2 (ref. \cite{Novikov1988})& 2 (ref. \cite{Nachman1988})& ? & 4 (this work)\\
$\bu$  &2 (ref. \cite{Agaltsov2016})& 2 (ref.  \cite{Zotov2017}) & ? & 2 (ref.  \cite{Zotov2017})\\
$c$, $\rho$, $\gamma$, $\bu$  &3 (ref. \cite{Agaltsov2016})& $\ge 3$ (ref. \cite{Shurup-priv})  & ? & ? \\
\hline 
\end{tabular}
\end{table}

{\bf Outstanding problems.}  
Further advances in local helioseismology will require improved forward solvers for vector MHD wave equations (see refs.~\cite{Cameron2008, Hanasoge2010}) and homogenized wave equations \cite{Hanasoge2013}, as well as improved inversion methods that minimize the number of forward solves \cite{Hanasoge2011,ref2}. A major challenge in local helioseismology is the very large size of the input dataset, e.g. $\sim 10^{12}$ pairs of points times $\sim 10^2$ frequencies in time-distance helioseismology. As a result, it is important to either select or average the input data before inverting them. One interesting averaging scheme that deserves further attention is helioseismic holography \cite{Lindsey2000a, Lindsey2000b, Skartlien2002}, which uses  Green's second identity to image scatterers in the Sun, as in Porter-Bojarski holography \cite{Porter1982,Devaney1985}.

\end{talk}

\end{document}